OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m =

OpenStudy (anonymous):

given slope ?

OpenStudy (anonymous):

m= 2/5

OpenStudy (anonymous):

(y+2)=2/5(x-5)

OpenStudy (anonymous):

We want it in standard form

OpenStudy (anonymous):

y+2=2/5x-2 2/5x - y=4